OpenStudy (anonymous):

solve −4(x + 1) − 3 = −3(x − 4).

pooja195 (pooja195):

−4(x+1)−3=−3(x−4) (−4)(x)+(−4)(1)+−3=(−3)(x)+(−3)(−4)(Distribute) −4x+−4+−3=−3x+12 (−4x)+(−4+−3)=−3x+12(Combine Like Terms) −4x+−7=−3x+12 −4x−7=−3x+12 Add 3x to both sides. −4x−7+3x=−3x+12+3x −x−7=12 Add 7 to both sides. −x−7+7=12+7 −x=19 Divide both dies by -1 to get answer

OpenStudy (johnn.):

Let's solve your equation step-by-step. −4(x+1)−3=−3(x−4) Step 1: Simplify both sides of the equation. −4(x+1)−3=−3(x−4) Simplify: −4x−7=−3x+12 Step 2: Add 3x to both sides. −4x−7+3x=−3x+12+3x −x−7=12 Step 3: Add 7 to both sides. −x−7+7=12+7 −x=19 Step 4: Divide both sides by -1. −x−1=19−1 x=−19 Answer: x=−19